m over 8 = 9 over five

what does m equal?

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

m over 8 = 9 over five

what does m equal?
==================
m/8 = 9/5
Cross multiply
5m = 72
m = 72/5 = 14.4

Answer by Theo(13342)

m/8 = 9/5

cross multiply to get 5m = 72

divide both sides of this equation by 5 to get m = 72/5.

if this solution is good, then (72/5) / 8 = 9/5

(72/5) / 8 is the same as 72 / (5*8) which is the same as 72 / 40

since 8 goes into 72 9 times and 8 goes into 40 5 times, then 72 / 40 is equivalent to 9 / 5

what you just did was divide both numerator and denominator by 8 which simplified the expression of 72/8 to make it equal to 9/5.

the result is that the equation (72/5) / 8 = 9/5 becomes 9/5 = 9/5 which is true.

this confirms that m = 72/5 is a solution to the equation.